# Piecewise M-Stationarity and Related Algorithms for Mathematical Programs with Complementarity Constraints

### 24 Mar 2026 v1

@@ 3.4 C-Stationarity Convergence
As $\epsilon^{k}$ tends to zero, the function $\Phi_{i}^{0}$ is in general not differentiable for $i\in\beta(\bar{z})$. However, if $\Phi_{i}^{0}(z)$ is locally Lipschitz [5, Section 1.2] near $\bar{z}$, the generalized gradient $\partial\Phi_{i}^{0}(\bar{z})$ is generated by a convex hull ([5, Theorem 2.5.1] [6, Eq.(3.1.5)])
@@
For $\delta_{i}\in\partial\Phi_{i}^{0}(\bar{z})$, it follows that ([26, Lemma 1])
@@ 4.1 Unbounded NLP Multipliers and Inaccurate Solution
with $v^{g,k},v^{G,k},v^{H,k},v^{REG,k}\geq 0$, the relations between the NLP multipliers $v^{k}=(v^{g,k},v^{h,k},v^{G,k},v^{H,k},v^{REG,k})$ at $z^{k}$ and the MPCC multipliers $\bar{\lambda}=(\bar{\lambda}^{g},\bar{\lambda}^{h},\bar{\lambda}^{G},\bar{\lambda}^{H})$ at $\bar{z}$ can be expressed by (see also [27, Eq.(6) and Theorem 3.1])
@@
It has been proved that $\bar{z}$ is S-stationary for MPCC (1) if and only if it is a stationary point of REG(0) [10, Proposition 4.1], provided that bounded multipliers exist.

### 14 Apr 2026 v2

@@ 3.4 C-Stationarity Convergence
As $\epsilon^{k}$ tends to zero, the function $\Phi_{i}^{0}$ is in general not differentiable for $i\in\beta(\bar{z})$. However, if $\Phi_{i}^{0}(z)$ is locally Lipschitz [4, Section 1.2] near $\bar{z}$, the generalized gradient $\partial\Phi_{i}^{0}(\bar{z})$ is generated by a convex hull ([4, Theorem 2.5.1] [5, Eq.(3.1.5)])
@@
For $\delta_{i}\in\partial\Phi_{i}^{0}(\bar{z})$, it follows that ([25, Lemma 1])
@@ 3.5 M-stationarity Convergence
It is worth noting that this theorem assumes MPCC-MFCQ together with MPCC-CRCQ, which is still weaker than MPCC-LICQ. To see this, consider the following constraints: $z_{1}+z_{3}\leq 0,z_{2}+z_{3}\leq 0,0\leq z_{1}\perp z_{2}\geq 0$, at point $z=(0,0,0)$.
@@ 4.1 Unbounded NLP Multipliers and Inaccurate Solution
with $v^{g,k},v^{G,k},v^{H,k},v^{REG,k}\geq 0$, the relations between the NLP multipliers $v^{k}=(v^{g,k},v^{h,k},v^{G,k},v^{H,k},v^{REG,k})$ at $z^{k}$ and the MPCC multipliers $\bar{\lambda}=(\bar{\lambda}^{g},\bar{\lambda}^{h},\bar{\lambda}^{G},\bar{\lambda}^{H})$ at $\bar{z}$ can be expressed by (see also [26, Eq.(6) and Theorem 3.1])
@@
It has been proved that $\bar{z}$ is S-stationary for MPCC (1) if and only if it is a stationary point of REG(0) [9, Proposition 4.1], provided that bounded multipliers exist.